OpenStudy (anonymous):

A system of equations is shown below. -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

looks like you just asked this right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. ok lets multiply the second equation by \(2\)

OpenStudy (anonymous):

\[-3x + 7y = -16 \\-9x + 5y = 16\] becomes \[-3x + 7y = -16 \\-18x + 10y = 32\] then add and get \[-21x+17y=16\]

OpenStudy (anonymous):

now your new system is \[-3x+7y=-16\\ -21x+17y=16\]

OpenStudy (anonymous):

that is my six points worth i guess your last job is to solve both systems of equations and see that you get the same answer

OpenStudy (anonymous):

ok